# Computation Mechanism Behind LLM Position Generalization

### v2

@@ 2.1 Self-Attention and Positional Encoding
Self-attention is the core design in most modern LLMs for information flow to words from their contexts Vaswani et al. (2017). It is also the primary (and often the only) component to inject text position information since the introduction of relative position encoding Su et al. (2021); Touvron et al. (2023); Dubey et al. (2024); OpenAI (2023), which is the subject of investigation in this work. Despite architecture variants, it is generally designed as a Softmax-based weighted average over contextual “value” vectors $\{{\bm{v}}_{j}|j\leq i\}$ before current position $i$. The average weights $w({\bm{q}}_{i},{\bm{k}}_{j},i-j)$ are determined by the relevance between the current word’s “query” vector ${\bm{q}}_{i}$, contextual words’ “key” vectors $\{{\bm{k}}_{j}|j\leq i\}$, and their relative position $i-j$. The output feature vector for the current token $\bm{o}_{i}$ is therefore:
@@
In spite of the existence of other choices of function $w(\cdots)$ like Alibi Press et al. (2021), the de-facto mainstream choice is RoPE Su et al. (2021). It decomposes ${\bm{q}}$ and ${\bm{k}}$ vectors into 2-D tuples and lets them rotate in angle $(i-j)\theta_{r}$, where each 2-D tuple $r$ has a different rotating “angular speed” $\theta_{r}$.
@@ 2.2 Position Generalization (of both LLMs and Humans)
which encodes pragmatic information such as emphasis Payne (1992). This aspect has been computationally measured (Kahane et al., 2023) and used to evaluate linguistic complexity Szmrecsanyi (2016).
@@ 5 Conclusions and Future Work
  • Improved efficiency in attention calculation: after estimating the semantic importance and positional patterns, some (regions of) attention logits can be estimated without calculating precisely. This will be especially useful when those logits are low so that estimation errors will only translate to a small deviation in their exponentials, which are used in the softmax operator later. Such techniques can be further incorporated into novel architectures, which will be inherently more efficient.
@@ Acknowledgement
This research is supported by U.S. IARPA BENGAL, DARPA INCAS Program No. HR001121C0165 and No. HR00112430325, and DARPA ITM Program No. FA8650-23-C-7316. The views, opinions, and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.

### v3

@@ 2.1 Self-Attention and Positional Encoding
Self-attention is the core design in most modern LLMs for information flow to words from their contexts Vaswani et al. (2017). It is also the primary (and often the only) component to inject text position information since the introduction of relative position encoding Su et al. (2021); Touvron et al. (2023); Dubey et al. (2024); OpenAI (2023), which is the subject of investigation in this work. Despite architecture variants, it is generally designed as a Softmax-based weighted average over contextual “value” vectors $\{{\bm{v}}_{j}|j\leq i\}$ before the current position $i$. The average weights $w({\bm{q}}_{i},{\bm{k}}_{j},i-j)$ are determined by the relevance between the current word’s “query” vector ${\bm{q}}_{i}$, contextual words’ “key” vectors $\{{\bm{k}}_{j}|j\leq i\}$, and their relative position $i-j$. The output feature vector for the current token $\bm{o}_{i}$ is therefore:
@@
Despite the existence of other choices of the function $w(\cdots)$ like Alibi Press et al. (2021), the de-facto mainstream choice is RoPE Su et al. (2021). It decomposes ${\bm{q}}$ and ${\bm{k}}$ vectors into 2-D tuples and lets them rotate in angle $(i-j)\theta_{r}$, where each 2-D tuple $r$ has a different rotating “angular speed” $\theta_{r}$.
@@ 2.2 Position Generalization (of both LLMs and Humans)
which encodes pragmatic information such as emphasis Payne (1992). This aspect has been computationally measured Kahane et al. (2023) and used to evaluate linguistic complexity Szmrecsanyi (2016).
@@ 5 Conclusions and Future Work
  • Improved attention calculation: after estimating the semantic importance and positional patterns, some (regions of) attention logits can be estimated without calculating precisely. This will be especially useful when those logits are low so that estimation errors will only translate to a small deviation in their exponentials. The insights offered in this work can also be used to inspire length extension methods of LMs, e.g., the explorations in later work Bianchessi et al. (2025).
@@ Acknowledgement
This research is supported by U.S. DARPA INCAS Program No. HR001121C0165, and DARPA ITM Program No. FA8650-23-C-7316. The views, opinions, and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.